Karen wants to advertise how many chocolate chips are in each Big Chip cookie at her bakery. She randomly selects a sample of help!
Answer #1
sample mean 'x̄= 15.800
sample size    n= 78.00
std deviation s= 1.900
std error ='sx=s/√n= 0.2151
for 95% CI; and 77 df, value of t= 1.9910
margin of error E=t*std error    = 0.428
lower bound=sample mean-E = 15.372
Upper bound=sample mean+E = 16.228
from above 95% confidence interval for population mean =(15.4 <μ< 16.2)